Sahar is analyzing the probability to open her own beauty salon. If she chooses to operate her own salon she would have to pay Rs. 70,000 p.m as rent and staff that would cost her Rs. 30,000 p.m and is expected to earn Rs. 150,000 p.m. While if she continues working as the head makeup artist in one of renowned salons of Karachi she would earn 120,000. Analyze the information above and find explicit and implicit cost.
1
Expert's answer
2020-11-12T17:48:50-0500

Explicit costs are normal business costs that directly affect the profitability of a business. In this case explicit cost are rent and wages. Therefore explicit cost is as shown below:

"Explicit Costs = Rent + wages\n = 70,000 +300000\n =100,000\n =Rs 100,000"


Implicit costs on the other hand are costs not clear defined or identified. They often deal with intangibles and usually described as opportunity costs. The implicit cost are as shown below: "Implicit Costs =120,000 - (150,000-100,000)\n = 120,000-50,0000\n =70,0000\n = Rs 70,000"